# {category}

### v5

@@ 1 Introduction
The e-value was specially designed to assess the logical truth value (a.k.a. statistical significance) of sharp (a.k.a. precise) hypotheses in the context of Bayesian statistics. The e-value has desirable asymptotic, geometrical (i.e., invariance), and logical (i.e., compositional) properties that allow consistent and coherent evaluation and testing of sharp statistical hypotheses. Furthermore, in applied modeling, the FBST offers an easy to implement and powerful statistical test that is fully compliant with Bayesian principles of good inference, like the likelihood principle, see [17, 91, 160, 248].
@@
In the context of statistical test of hypotheses, a compositional logic is conveyed by an algebraic formalism that allows the evaluation of truth-functions of composite models and truth-values of composite hypotheses by algebraic operations on the corresponding truth functions of elementary models and truth values of elementary hypotheses. The e-value and the FBST have a rich, expressive and intuitive compositional logic, while traditional truth-values and accompanying tests offered by either frequentist (classical) statistics, like the p-value, or by Bayesian statistics, like Bayes factors, have important and well-known deficiencies in this regard, specially in cases involving sharp statistical hypotheses. Furthermore, logically coherent evaluations and testing of sets and sub-sets of statistical hypotheses should render sequences of inferential reasoning that do not generate internal contradictions or anti-intuitive results. As expected, the e-value and the FBST comply with well-established rules of logical coherence, even in the case of sharp hypotheses, while traditional alternatives often fail to do so.
@@
Asymptotically, $\mbox{sev}(H|X)$ – the standardized e-value – shares several properties of the p-value, the well-known significance measure of frequentist statistics. This allows the use of the standardized e-value in frequentist oriented applications, retaining, nevertheless, many theoretical characteristics of the Bayesian framework. Many theoretical developments and practical applications of the e-value and the FBST have already been published in the scientific literature, see [169]. Concise entries about the e-value and the FBST are available at the International Encyclopedia of Statistical Science and online at Wiley’s StatsRef, see [55, 167].
@@
Section 2 reviews the Bayesian statistical framework; Section 3 defines the e-value; Sections 4, 5, and 6 explain the invariance, asymptotic and compositional properties of the e-value; Section 7 defines the GFBST - the Generalized Full Bayesian Significance Test and its logical properties; Sections 8 and 9 comment on computational implementation and give a detailed numerical example in model selection; Section 10 lists a representative assortment of articles from many practical applications of the e-value and the FBST already published in the scientific literature. Section 11 considers the philosophical consequences of the aforementioned developments by briefly commenting on the Objective cognitive constructivism epistemological framework, which was specifically developed to accommodate the formal properties of the e-value and the FBST, and renders a naturalized approach to ontology and metaphysics. Section 12 presents some topics for further research at the interface between Logic and Statistics. One of the objectives of this paper is to foster work in areas of interface or overlap between Logic and Statistics and to stimulate greater cooperation between the two communities. Hence, for the sake of clarity and completeness, this paper includes some basic definitions and explanations in both areas that may be convenient to facilitate mutual understanding of more advanced topics.
@@ 3 The e-value
In the special case of the flat reference density, $r(\theta)\propto 1$, the surprise function coincides with the posterior probability density function. Its upper cuts, $\overline{T}(v)$, coincide with highest probability density sets (HPDS), making it easier to visualize all the elements defined in the last paragraph. Figure 1 depicts a simple example with a unimodal surprise function and a simply connected tangential set, see [161, Sec.4.3]. However, in general, the posterior probability density function may have multiple local maxima, the tangential set may have multiple connected components, and all these components must be taken into account when computing the e-value. In this respect, the e-value differs from traditional uses of HPDS that only consider a single connected component, usually the connected component containing $\widehat{\theta}$, the unconstrained maximum a posteriory estimator, see [28, 57, 139, 194, 224, 249].
@@ 5 Asymptotic Consistency
(b) If $H$ is true, i.e. $\theta^{0}\in H$; and $H$ is sharp, i.e. $\dim(H)<\dim(\Theta)$; Then $\mbox{sev}(H\,|\,X)$ converges (in distribution) to $U[0,1]$, the uniform distribution in the unit interval. The standardized e-value shares these asymptotic properties with the p-value – the well-known and widely used significance measure of frequentist statistics. Hence, in a frequentist oriented application, one could use the e-value as a Bayesian replacement for the p-value, preserving already familiar decision procedures and their interpretations. Moreover, in several modeling applications, the e-value exhibits better convergence characteristics than those of the p-value, see [19, 131, 134, 132, 133]. Further consistency properties of the e-value and higher order asymptotic approximations have been studied and developed in [33, 172, 175, 189, 190, 237, 238, 239].
@@ 8 Computational Implementation
– computing an approximation of the truth function $W(v)$. In most cases, this task can be accomplished by using a computational condensation procedure, like [118, 251], on a numerical sequence, $s^{(k)}=s(\theta^{(k)})$, obtained by sampling from the posterior distribution, $p_{n}(\theta)$. Sampling sequences can, in turn, be generated using Markov Chain Monte-Carlo methods, see [31, 92, 98], or variations thereof, like the Hit-and-Run algorithm, see [121];
@@ 9 Numerical Example in Model Selection
Figure 7 (left) depicts the polynomial fitting problem for the classical Sakamoto et al. benchmark dataset presented in [193, ch.8]. This dataset, given in Table 1, was generated by a simulation from the i.i.d. (independent and identically distributed) stochastic process
@@
Figure 7: Benchmark (left) and alternative (right) data points, $\circ$, for Sakamoto et al. [193, ch.8] polynomial fitting problem; Exponential target function, $\diamond$; Best fitted polynomials, — , of order 0 to 4; and Order 2 polynomial, ***, that renders the smallest regularized error for both datasets.
@@
Order $R_{EMP}$ $R_{FPE}$ $R_{SBC}$ $R_{GCV}$ $R_{SMS}$ $R_{AIC}$ $\,\mbox{ev}\,(H)$ $\mbox{sev}(H)$
0 0.03712 0.04494 0.04307 0.04535 0.04419 -07.25 0.000 0.000
1 0.02223 0.02964 0.02787 0.03025 0.02858 -20.35 0.009 0.001
2 0.01130 0.01661 0.01534 0.01724 0.01560 -32.13 0.013 0.000
3 0.01129 0.01835 0.01667 0.01946 0.01667 -30.80 0.999 0.818
4 0.01088 0.01959 0.01751 0.02133 0.01710 -29.79 0.995 0.403
5 0.01087 0.02173 0.01913 0.02445 0.01811 -27.86 0.999 0.641
Table 2: Sakamoto et al. [193, ch.8] benchmark dataset for the polynomial fitting problem
@@ 11 Epistemology, Ontology, and Metaphysics
Historically, the development of statistical significance measures (or logical truth-values), and the corresponding tests of hypotheses used in statistical science, have been reciprocally influenced by the epistemological frameworks in which they are presented. Rev. Thomas Bayes (1701-1761), whose work was communicated posthumously by Rev. Richard Price (1723-1791), developed the first methods in this area with clear goals in mind, as stated in [16], also quoted in [11, p.84] and [229, p.245]:
@@ 12 Future Research and Final Remarks
In December 2018, at CLE-UNICAMP, Walter Carnielli organized the workshop Induction, Probability and their Dilemmas, where the first author gave the presentation The Problem of Induction in Statistical Science. During this workshop, we discussed several topics for further research, most of them motivated by areas of common interest that, nevertheless, are usually approached quite differently by the communities of Logic and Statistics. This section presents four of these topics.
@@
In Luhmann’s theory of law, the main purpose of the legal system is congruent generalization of normative behavior expectations, see [144, pp.77,82]. This emblematic statement can be interpreted as follows, see [228]: In Luhmann’s view, norms are neither preexisting conditions nor a priori factual realities. Instead, norms are conceived as intentional projects or idealized models of how society should be; see [144, p.40]. Moreover, in such idealized models, social harmony is based on establishing well-defined, stable, sustainable, and reliable behavioral patterns, also known as eigen-behaviors or eigen-solutions, see [79, 228]. Furthermore, laws and regulations of a society should reflect its norms, stimulating/ penalizing forms of conduct that sustain/ disrupt virtuous eigen-behaviors. Finally, in Luhmann’s view, social norms are not static but essentially dynamic, co-evolving (over long runs) with the behavioral patterns in the society that they simultaneously try to describe, regulate and stabilize (in their present forms, over short runs). In this context, we can try to reinterpret some lessons about model selection and their complexity learned in Section 9, claiming that, as in the case of good empirical laws, good social laws should follow the golden path of equilibrium, avoiding extremes of scarcity and excess, see also [216, 221].
@@ Acknowledgments
The authors take this opportunity to thank Prof. Walter Carnielli for, along the years, inviting, welcoming, and stimulating investigations, discussions, and interactions at the borders or areas of overlap between logic, probability, statistics, formal methods in science, philosophy of science and general philosophy. The authors have benefited greatly from Walter Carnielli’s encyclopedic knowledge, open-mindedness, positive attitude, and active engagement in fostering research in all these areas, either at his home base at CLE - the Center for Logic, Epistemology and History of Science of UNICAMP, or throughout Brazil and around the world. The authors also thank Profs. Abilio Rodrigues, Alfredo Freire and Henrique Antunes for editing this special volume celebrating Walter’s 70th birthday.
@@
The authors have received direct or indirect support from the following institutions: USP – University of São Paulo (JMS, CABP, MSL, LGE, and WSB); UFSCar – Federal Univesity of São Carlos (RI, RBS); ABJ – Brazilian Jurimetrics Association (JMS, RBS); CNPq – Brazilian National Counsel of Technological and Scientifc Development, grants PQ 303290/2021-8 (JMS), PQ 302767/2017-7 (CABP), and PQ 309607-2020-5 (RI); FAPESP – State of São Paulo Research Foundation, grants 2019/11321-9 (RI), CEPID CeMEAI – Center for Mathematical Sciences Applied to Industry 2013/07375-0 (JMS, MSL), and CEPID Shell-RCGI – Research Center for Greenhouse Gas Innovation 2014/50279-4 (JMS); and CAPES – Coordination for Improvement of Higher Education Personnel, finance code 001 (RI, RBS, MAD). Concerning this article, the authors have no conflicts of interest to declare.

### v6

@@ 1 Introduction
The e-value was specially designed to assess the logical truth value (a.k.a. statistical significance) of sharp (a.k.a. precise) hypotheses in the context of Bayesian statistics. The e-value has desirable asymptotic, geometrical (i.e., invariance), and logical (i.e., compositional) properties that allow consistent and coherent evaluation and testing of sharp statistical hypotheses. Furthermore, in applied modeling, the FBST offers an easy-to-implement and powerful statistical test that is fully compliant with Bayesian principles of good inference, like the likelihood principle, see [17, 95, 165, 257].
@@
In the context of statistical test of hypotheses, a compositional logic is conveyed by an algebraic formalism that allows the evaluation of truth-functions of composite models and truth-values of composite hypotheses by algebraic operations on the corresponding truth functions of elementary models and truth values of elementary hypotheses. The e-value and the FBST have a rich, expressive, and intuitive compositional logic, while traditional truth-values and accompanying tests offered by either frequentist (classical) statistics, like the p-value, or by Bayesian statistics, like Bayes factors, have important and well-known deficiencies in this regard, specially in cases involving sharp statistical hypotheses. Furthermore, logically coherent evaluations and testing of sets and sub-sets of statistical hypotheses should render sequences of inferential reasoning that do not generate internal contradictions or anti-intuitive results. As expected, the e-value and the FBST comply with well-established rules of logical coherence, even in the case of sharp hypotheses, while traditional alternatives often fail to do so.
@@
Asymptotically, $\mbox{sev}(H|X)$ – the standardized e-value – shares several properties of the p-value, the well-known significance measure of frequentist statistics. This allows the use of the standardized e-value in frequentist-oriented applications, retaining, nevertheless, many theoretical characteristics of the Bayesian framework. Many theoretical developments and practical applications of the e-value and the FBST have already been published in the scientific literature, see [174]. Concise entries about the e-value and the FBST are available at the International Encyclopedia of Statistical Science and online at Wiley’s StatsRef, see [58, 172].
@@
Section 2 reviews the Bayesian statistical framework; Section 3 defines the e-value; Sections 4, 5, and 6 explain the invariance, asymptotic, and compositional properties of the e-value; Section 7 defines the GFBST - the Generalized Full Bayesian Significance Test and its logical properties; Sections 8 and 9 comment on computational implementation and give a detailed numerical example in model selection; Section 10 lists a representative assortment of articles from many practical applications of the e-value and the FBST already published in the scientific literature. Section 11 considers the philosophical consequences of the aforementioned developments by briefly commenting on the Objective cognitive constructivism epistemological framework, which was specifically developed to accommodate the formal properties of the e-value and the FBST, and renders a naturalized approach to ontology and metaphysics. Section 12 presents some topics for further research at the interface between Logic and Statistics. One of the objectives of this paper is to foster work in areas of interface or overlap between Logic and Statistics and to stimulate greater cooperation between the two communities. Hence, for the sake of clarity and completeness, this paper includes some basic definitions and explanations in both areas that may be convenient to facilitate mutual understanding of more advanced topics.
@@ 3 The e-value
In the special case of the flat reference density, $r(\theta)\propto 1$, the surprise function coincides with the posterior probability density function. Its upper cuts, $\overline{T}(v)$, coincide with highest probability density sets (HPDS), making it easier to visualize all the elements defined in the last paragraph. Figure 1 depicts a simple example with a unimodal surprise function and a simply connected tangential set, see [166, Sec.4.3]. However, in general, the posterior probability density function may have multiple local maxima, the tangential set may have multiple connected components, and all these components must be taken into account when computing the e-value. In this respect, the e-value differs from traditional uses of HPDS that only consider a single connected component, usually the connected component containing $\widehat{\theta}$, the unconstrained maximum a posteriory estimator, see [28, 61, 143, 199, 233, 258].
@@ 5 Asymptotic Consistency
(b) If $H$ is true, i.e. $\theta^{0}\in H$; and $H$ is sharp, i.e. $\dim(H)<\dim(\Theta)$; Then $\mbox{sev}(H\,|\,X)$ converges (in distribution) to $U[0,1]$, the uniform distribution in the unit interval. The standardized e-value shares these asymptotic properties with the p-value – the well-known and widely used significance measure of frequentist statistics. Hence, in a frequentist-oriented application, one could use the e-value as a Bayesian replacement for the p-value, preserving already familiar decision procedures and their interpretations. Moreover, in several modeling applications, the e-value exhibits better convergence characteristics than those of the p-value, see [19, 135, 138, 136, 137]. Further consistency properties of the e-value and higher order asymptotic approximations have been studied and developed in [33, 177, 180, 194, 195, 246, 247, 248].
@@ 8 Computational Implementation
– computing an approximation of the truth function $W(v)$. In most cases, this task can be accomplished by using a computational condensation procedure, like [122, 260], on a numerical sequence, $s^{(k)}=s(\theta^{(k)})$, obtained by sampling from the posterior distribution, $p_{n}(\theta)$. Sampling sequences can, in turn, be generated using Markov Chain Monte-Carlo methods, see [31, 96, 102], or variations thereof, like the Hit-and-Run algorithm, see [44, 125];
@@ 9 Numerical Example in Model Selection
Figure 7 (left) depicts the polynomial fitting problem for the classical Sakamoto et al. benchmark dataset presented in [198, ch.8]. This dataset, given in Table 1, was generated by a simulation from the i.i.d. (independent and identically distributed) stochastic process
@@
Figure 7: Benchmark (left) and alternative (right) data points, $\circ$, for Sakamoto et al. [198, ch.8] polynomial fitting problem; Exponential target function, $\diamond$; Best fitted polynomials, — , of order 0 to 4; and Order 2 polynomial, *, that renders the smallest regularized error for both datasets.
@@
Order $R_{EMP}$ $R_{FPE}$ $R_{SBC}$ $R_{GCV}$ $R_{SMS}$ $R_{AIC}$ $\,\mbox{ev}\,(H)$ $\mbox{sev}(H)$ 0 0.03712 0.04494 0.04307 0.04535 0.04419 -07.25 0.000 0.000 1 0.02223 0.02964 0.02787 0.03025 0.02858 -20.35 0.009 0.001 2 0.01130 0.01661 0.01534 0.01724 0.01560 -32.13 0.013 0.000 3 0.01129 0.01835 0.01667 0.01946 0.01667 -30.80 0.999 0.818 4 0.01088 0.01959 0.01751 0.02133 0.01710 -29.79 0.995 0.403 5 0.01087 0.02173 0.01913 0.02445 0.01811 -27.86 0.999 0.641
Table 2: Sakamoto et al. [198, ch.8] benchmark dataset for the polynomial fitting problem
@@ 11 Epistemology, Ontology, and Metaphysics
Historically, the development of statistical significance measures (or logical truth-values), and the corresponding tests of hypotheses used in statistical science, have been reciprocally influenced by the epistemological frameworks in which they are presented. Rev. Thomas Bayes (1701-1761), whose work was communicated posthumously by Rev. Richard Price (1723-1791), developed the first methods in this area with clear goals in mind, as stated in [16], also quoted in [11, p.84] and [238, p.245]:
@@ 12 Future Research and Final Remarks
In December 2018, at CLE-UNICAMP, Walter Carnielli organized the workshop Induction, Probability and their Dilemmas, where the first author gave the presentation The Problem of Induction in Statistical Science. During this workshop, we discussed several topics for further research, most of them motivated by areas of common interest that, nevertheless, are usually approached quite differently by the communities of Logic and Statistics. Four of these topics are presented in the sequel, followed by some additional topics for further research.
@@
In Luhmann’s theory of law, the main purpose of the legal system is congruent generalization of normative behavior expectations, see [149, pp.77,82]. This emblematic statement can be interpreted as follows, see [237]: In Luhmann’s view, norms are neither preexisting conditions nor a priori factual realities. Instead, norms are conceived as intentional projects or idealized models of how society should be; see [149, p.40]. Moreover, in such idealized models, social harmony is based on establishing well-defined, stable, sustainable, and reliable behavioral patterns, also known as eigen-behaviors or eigen-solutions, see [83, 237]. Furthermore, laws and regulations of a society should reflect its norms, stimulating/ penalizing forms of conduct that sustain/ disrupt virtuous eigen-behaviors. Finally, in Luhmann’s view, social norms are not static but essentially dynamic, co-evolving (over long runs) with the behavioral patterns in the society that they simultaneously try to describe, regulate and stabilize (in their present forms, over short runs). In this context, we can try to reinterpret some lessons about model selection and their complexity learned in Section 9, claiming that, as in the case of good empirical laws, good social laws should follow the golden path of equilibrium, avoiding extremes of scarcity and excess, see also [225, 230].
@@
Individual vs. Collective Liability, Legal Burden of Proof, etc.
As commented at the end of Section 2, the Onus Probandi and In Dubio Pro Reo principles, as they are used in the legal system, were a source of inspiration for the definitions of the e-value and the FBST, see [94, 99, 133, 167, 168, 176, 222, 237]. Moreover, it has long been recognized that many decision procedures traditionally used in Bayesian statistics are incompatible with these principles. Dennis Lindley uses the famous gatecrasher example to argue that, under certain circumstances, collective responsibility is a preferable principle of justice, and that legal decision procedures implying collective liabilities and punishments should be accepted, see [144, 47]. Furthermore, Lindley shows how legal decisions based on Bayes Factors are compatible with such a collective liability principle.
The right tool for the right job! In accordance with this dictum, further research could contrast the legal and social implications of alternative liability principles, their potential or desirability, and their compatibility with alternative inference theories, decision procedures, and epistemological frameworks.
e-value and FBST under Non-Standard Regularity Conditions
The asymptotic consistency and standardization procedures presented in Section 5 rely on standard regularity conditions that include continuity and differentiability of several functions defining the statistical model, and also the location of the true value of the parameter at the topological interior of the parameter space. However, many interesting models present non-standard conditions. For example, mixture models and separate hypotheses models often present hypotheses at the (topological) border of the parameter space, see [136, 138], whereas eigenvalue and eigenvector inference problems may present far more challenging conditions, see [202]. The methods presented in [46, 59, 205] and references therein provide tools to generalize the e-value standardization procedure and the corresponding asymptotic results to non-standard regularity conditions.
Other non-standard inference conditions arise in semi-parametric models like, for example, Hilbert space models based on Fourier, orthogonal polynomials, wavelets, and other infinite functional bases. In these cases, it is often necessary to use weakly informative priors that dampen high-frequency modes that effectively reduce the size of resulting truncated models, see [240, Sec.6] and references therein. Further research should explore convenient forms and characteristics of such weakly informative priors and related topics.
Efficient Computational Methods and Implementations
The e-value and FBST research program would greatly benefit from computational tools tailor-made for computational efficiency, reliability, and ease of use in the numerical tasks described at Section 8. Most beneficial would be the construction of an open-source and user-friendly environment combining specially adapted versions of the following techniques: (1) Efficient Monte-Carlo integrators based on Hit-and-Run, Nested Sampling and similar algorithms, see [44, 125, 214, 215]; (2) Efficient computational condensation procedures used in conjunction with the aforementioned Monte-Carlo procedures, see [122, 260]; (3) Efficient stochastic optimization methods working in conjunction with the aforementioned Monte-Carlo procedures, see [175, 216, 220, 251, 253]; (4) Higher order asymptotic approximations used either to develop stand-alone fast e-value calculation procedures or to develop variance reduction techniques for the aforementioned Monte-Carlo procedures, see [33, 177, 180, 194, 195, 246, 247, 248]; (5) Efficient numerical convolution procedures, used in conjunction with the aforementioned condensation procedures, as required in the analysis of complex models as described in Section 6.
@@ Acknowledgments
The authors take this opportunity to thank Prof. Walter Carnielli for, along the years, inviting, welcoming, and stimulating investigations, discussions, and interactions at areas of overlap between logic, probability, statistics, formal methods in science, philosophy of science, and general philosophy. The authors have benefited greatly from Walter Carnielli’s encyclopedic knowledge, open-mindedness, positive attitude, and active engagement in fostering research in all these areas, either at his home base at CLE - the Center for Logic, Epistemology, and History of Science of UNICAMP, or throughout Brazil and around the world. The authors are grateful to Profs. Abilio Rodrigues, Alfredo Freire and Henrique Antunes for organizing a Festschrift celebrating Walter’s 70th birthday.
The authors thank the organizers of the following conferences, where this paper was presented: CLMPST-2023, the 17th Congress on Logic, Methodology and Philosophy of Science and Technology, at FCE-UBA, Buenos Aires, July 24-29, 2023; Principia-2023, the 13th Principia International Symposium, at UFSC, Florianópolis, August 14-17, 2023; RatioLog-2024, the 2nd Workshop on Logic, Rationality, and Probability, at CLE-Unicamp, Campinas, April 22-25, 2024; and MaxEnt-2024, the 43rd International Workshop on Bayesian Inference and Maximum Entropy Methods in Science and Engineering, at U-Ghent, July 1-5, 2024.
@@
The authors have received direct or indirect support from the following institutions: USP – University of São Paulo (JMS, CABP, MSL, LGE, and WSB); UFSCar – Federal Univesity of São Carlos (RI, RBS); ABJ – Brazilian Jurimetrics Association (JMS, RBS); CNPq – Brazilian National Counsel of Technological and Scientific Development, grants PQ 303290/2021-8 (JMS), PQ 302767/2017-7 (CABP), and PQ 309607-2020-5 (RI); FAPESP – State of São Paulo Research Foundation, grants 2019/11321-9 (RI), CEPID CeMEAI – Center for Mathematical Sciences Applied to Industry 2013/07375-0 (JMS, MSL), and CEPID Shell-RCGI – Research Center for Greenhouse Gas Innovation 2014/50279-4 (JMS); and CAPES – Coordination for Improvement of Higher Education Personnel, finance code 001 (RI, RBS, MAD). Concerning this article, the authors have no conflicts of interest to declare.